Wheeling Company is a merchandiser that provided a balance sheet as of September 30 as shown below:

Wheeling Company
Balance Sheet
September 30
Assets
Cash $ 59,000
Accounts receivable 90,000
Inventory 32,400
Buildings and equipment, net of depreciation 214,000
Total assets $ 395,400
Liabilities and Stockholders’ Equity
Accounts payable $ 73,000
Common stock 216,000
Retained earnings 106,400
Total liabilities and stockholders’ equity $ 395,400

The company is in the process of preparing a budget for October and has assembled the following data:

Sales are budgeted at $240,000 for October and $250,000 for November. Of these sales, 35% will be for cash; the remainder will be credit sales. Forty percent of a month’s credit sales are collected in the month the sales are made, and the remaining 60% is collected in the following month. All of the September 30 accounts receivable will be collected in October.

The budgeted cost of goods sold is always 45% of sales and the ending merchandise inventory is always 30% of the following month’s cost of goods sold.

All merchandise purchases are on account. Thirty percent of all purchases are paid for in the month of purchase and 70% are paid for in the following month. All of the September 30 accounts payable to suppliers will be paid during October.

Selling and administrative expenses for October are budgeted at $78,000, exclusive of depreciation. These expenses will be paid in cash. Depreciation is budgeted at $2,000 for the month.

Required:

1. Using the information provided, calculate or prepare the following:

a. The budgeted cash collections for October.

b. The budgeted merchandise purchases for October.

c. The budgeted cash disbursements for merchandise purchases for October.

d. The budgeted net operating income for October.

e. A budgeted balance sheet at October 31.

2. Assume the following changes to the underlying budgeting assumptions:

(1) 50% of a month’s credit sales are collected in the month the sales are made and the remaining 50% is collected in the following month, (2) the ending merchandise inventory is always 10% of the following month’s cost of goods sold, and (3) 20% of all purchases are paid for in the month of purchase and 80% are paid for in the following month. Using these new assumptions, calculate or prepare the following:

a. The budgeted cash collections for October.

b. The budgeted merchandise purchases for October.

c. The budgeted cash disbursements for merchandise purchases for October.

d. The budgeted net operating income for October.

e. A budgeted balance sheet at October 31.
Answer #1

1. a. cash collection budget

October cash sales [$240,000 * 35%]    = 84000

Collection on accounts:

October sales[($240,000 * 65%) * 40%] = 62400

September Accounts receivable =90,000

Total cash collected =236400

b budgeted merchandise purchases

      October

sales $240,000   

Cost of goods sold {45% of sale} 108000   

add: Ending inventory[30%*112500]    33750   

Total material required 141750

less: Beginning inventory[108000*30%] 32400

   Budgeted purchase $109350

c. budgeted cash disbursements for merchandise purchases for October

October cash purchase [109350 * 30% ] =$32805

September Payable =73,000

   Total cash disbursements 105805

d. budgeted net operating income for October

      sales $240,000

less: cost of goods sold 108000   

Gross profit 132000

less: Depreciation 2000

Selling and administrative expenses $78,000

   Net income 52000

e. budgeted balance sheet at October 31

   Assets

   Cash [59,000 beginning +236400 - 105805 - $78,000] 111595

Accounts receivable [($240,000 * 65%) * 60 %] 93600

Inventory    33750  

Buildings and equipment, net of depreciation [214000 - 2000] 212000

  Total assets    450945

Liabilities and Stockholders’ Equity

   Accounts payable[109350 * 70% ] 76545

Common stock   216,000

Retained earnings [106,400 + 52000  ] 158400

Total liabilities and stockholders’ equity    450945

2   a.      cash collection budget

October cash sales [$240,000 * 35%]    = 84000

Collection on accounts:

October sales[($240,000 * 65%) * 50%] = 78000

September Accounts receivable =90,000

Total cash collected =252000

b     budgeted merchandise purchases

      October

sales $240,000   

Cost of goods sold {45% of sale} 108000   

add: Ending inventory[10%*112500]    11250   

Total material required 119250

less: Beginning inventory    32400

   Budgeted purchase $86850

c. budgeted cash disbursements for merchandise purchases for October

October cash purchase [86850 * 20% ] = 17370

September Payable =73,000

   Total cash disbursements =90370

d.    budgeted net operating income for October

      sales $240,000

less: cost of goods sold 108000

Gross profit 132000

less: Depreciation 2000

Selling and administrative expenses    $78,000

   Net income 52000

e.     budgeted balance sheet at October 31

     Assets

     Cash [59,000 beginning +252000 - 90370 - $78,000] 142630

Accounts receivable [($240,000 * 65%) * 50 %] 78000

Inventory    11250   

Buildings and equipment, net of depreciation [214000 - 2000] 212000

  Total assets 443880

Liabilities and Stockholders’ Equity

   Accounts payable 69480

Common stock 216,000

Retained earnings [106,400 + 52000 ] 158400

Total liabilities and stockholders’ equity    443880
Answer #2

Solution:

1(a) Amount $ 84,000 Budgeted Cash Collections for October Particulars Cash Sales (240,000 * 35%) Cash Collection in the mont1(e) Budgeted Balance Sheet for October ASSETS Cash Accounts Receivable Inventory Buildings and Equipment Amount $ 111,595 932 (d) Budgeted Net Operating Income for October Particulars Net Sales Cost of Goods Sold Gross Margin Selling & Administrativ